Find all values of K for which the quadratic equation has no real solutions.
8x^2+7x+k=0. Write your answer as an equality or inequality in terms of
k.
1
Expert's answer
2019-09-01T17:30:25-0400

Solution:

8x^2+7x+k=0

D=b^2-4ac=49-4*8*k

quadratic Equation has no real solution if D<0:

49-4*8*k<0

49<32k

Answer:

k>49/32